a man can swim at 6 km/hour. The speed of a stream is 2 km/hour. How much time will the man take to swim 400m if(a) he is swimming with the stream?(b) he is swimming against the stream?

Grade:10

1 Answers

Arun
6 years ago
Dear Student
 
(a)
if he is swimming with the stream, the resultant velocity = 6+2 = 8 km/hr
 = 8 * 5/18 = 20/9  m/sec
now
time  = distance /speed = 400/ (20/9)
 = 180 sec = 3 minutes
 
(b)
if against the stream, resultant velocity = 6 – 2 = 4 km/hr
 = 4 * 5/18 = 10/9 m/sec
hence time taken  = 400 / (10/9)
 = 360 sec = 6 minutes
 
Regards
Arun (askIITians forum expert)